Identify and graph the conic section.
4x=(y+4)^2-2

Answer by lwsshak3(11628)

Identify and graph the conic section.
4x=(y+4)^2-2
**
4x=(y+4)^2-2
4x+2=(y+4)^2
2(2x+1)=(y+4)^2
(y+4)^2=2(2x+1)
This is an equation for a parabola of the standard form: (y-k)^2=4p(x-h), with (h,k) being the (x,y) coordinates of the vertex. Parabola opens rightwards with a horizontal axis of symmetry.
..
For given equation:
Vertex: (-1/2, -4)
Axis of symmetry: y=-4
See graph below:
..
y=±(4x+2)^.5-4